OpenStudy (anonymous):

a softball has a volume of 126/6 pi cubic inches.find the radius of the softball

OpenStudy (anonymous):

i will show you to get redius

OpenStudy (anonymous):

great

OpenStudy (acxbox22):

volume=\[4/3\pi r ^{3}\]

OpenStudy (anonymous):

3.96is ans

OpenStudy (anonymous):

as we know that volume of sphere is 4/3 pi r^3 126/6*pi=4/3*pi*r^3

OpenStudy (anonymous):

solve for finding r